1. find the 7th term in 16, 8, 4,....

2. Find the sum of the first 11 terms in the sequence 4,6,9...

3. What is the sum of the first 100 even numbers

Answer by stanbon(75887)

1. find the 7th term in 16, 8, 4,....
a = 7 ; r = (1/2)
a(7) = 7*(1/2)^6 = 7/64
------------------------------------
2. Find the sum of the first 11 terms in the sequence 4,6,9...
a = 4 ; r = 3/2
S(11) = 4[(3/2)^11-1 / (3/2)-1]
-----
S(11) = 4[85.50/0.5]
S(11) = 4[171] = 684
-------------------------------------
3. What is the sum of the first 100 even numbers
---
2+4+6+....+200
200+198+....+2
-----
Adding you get
2S = 100(1st+last)
S = 50*202
Sum = 10,100
====================
Cheers,
Stan H.